OpenStudy (anonymous):

What is the solution of the system? Use substitution. y = x + 6 y = 2x

OpenStudy (anonymous):

its either y=x/7 or y=x+7

OpenStudy (anonymous):

A. (6, 12) B. (–12, –6) C. (–6, –12) D. (2, 4)

OpenStudy (anonymous):

nvm idk